OpenStudy (littleheron):

WILL MEDAL AND FAN Solve for y. -9x + 3y = 15 y = 9x + 5 y = -3x + 5 y = 3x + 5 y = 3x - 5

OpenStudy (abmon98):

3y=15+9x y=3(5+3x)/3 y=5+3x